OpenStudy (anonymous):

can someone factor x2 +x + 1

hero (hero):

complex?

OpenStudy (anonymous):

\[ \left(x+1/2-1/2\cdot\sqrt{3}\cdot i\right)\left(x+1/2+1/2\cdot \sqrt{3}\cdot i\right) \]

hero (hero):

That looks ugly

OpenStudy (anonymous):

Always looks ugly if you factor a polynomial with non-real zeros

OpenStudy (anonymous):

except \[ x^2+1=(x-i)(x+i) \] and the like :D

OpenStudy (anonymous):

thanks guys